OpenStudy (anonymous):

Find the value of each logarithm using the change of base formula. --> CLICK 2 SEE.

OpenStudy (anonymous):

\[LOG _{3}0.9375\]

OpenStudy (anonymous):

\[log_ab=\frac{log b}{log a}\]

OpenStudy (anonymous):

Remember the change of base formula is: \[Log_x a = \frac{ \log_{10} a }{ \log_{10} x }=\frac{ \log a }{ \log x }\]

OpenStudy (anonymous):

so \[ \frac{ \log 0.9375 }{\log 3 }\] ?

OpenStudy (anonymous):

Yes, just evaluate that in your calculator.

OpenStudy (anonymous):

but if my teacher would say find the exact value, would i leave it in this form? because we can't use calculators on the test.

OpenStudy (anonymous):

Yes, then you leave it like this.

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

No problem. Hope it helped.